What is the current share value of Rising Bank's stock to an investor who has a required rate of return of 25 percent? The current dividend is $5.45(D0) and the dividends are expected to grow 8 percent per year for 4 years. At the end of 4 years, the investor expects to sell the security for $50.

Question 14 options:

$44.60

$20.48

$32.77

$35.81
Answer #1

B14 fx Current Share Price = 35.81 [ Option D is the correct answer] А B F G 1 2 Years 3 4 Dividend Discount Rate @25% PresenB14 X fx Current Share Price = 35.81 [ Option D is the correct answer] A B 1 2 Years 3 1 4 2 Dividend =5.45*(1+8%) =C3*(1+8%)